OpenStudy (anonymous):

10x + 7y = 99 y − x = 2 Solve for the system shown. Put the answer in (x, y) format.

OpenStudy (anonymous):

10x + 7y = 99--------eq1 y − x = 2-------eq2 take eq 2 y-x=2 y=2+x put this value in eq1 10x+7(2+x)=99 10x+14+7x=99 17x=99-14 17x=85 x=85/17 x=5 now put the value of x in eq2 y-x=2 y-5=2 y=2+5 y=7 so x=5 and y=7 hope it helps :)